Given the following bill-of-materials: One A requires 2 units of B, 3 units of C, and 1 unit of D. One B requires 1 unit of C and 1 unit of D. One C requires 1 unit of D and 2 units of E. One D requires 1 unit of F and 2 units of G. If the demand for product A is 50 units, what will be the gross requirement for component E?
Answer #1

In order to produce 1 unit of A, number of C required is 3

Therefore in order to produce 50 units of A, the number of C required would be 3*50

= 150

In order to produce 1 Units of A, the number of B required is 2

Therefore in order to produce 50 units if A, the number of B required as 2*50

= 100

In order to produce 1 Unit of B, the number of C required is 1

Therefore in order to produce 100 units of B, the number of C required would be 1*100

= 100

Total C required would be 150+100

= 250

In order to produce 1 Unit of C, the number of E required is 2

Therefore, in order to produce 250 Units of C, the number of E required would be 2*250

= 500

In order to produce 50 Units of A, the number of E required would be 500
Answer #2

Given BOM Structure:

  1. 1 A requires:

    • 2 B

    • 3 C

    • 1 D

  2. 1 B requires:

    • 1 C

    • 1 D

  3. 1 C requires:

    • 1 D

    • 2 E

  4. 1 D requires:

    • 1 F

    • 2 G

(Note: E is only directly required by C.)



Step-by-Step Calculation:

1. Requirements for 50 A:

  • B: 50A×2B/A=100B

  • C: 50A×3C/A=150C

  • D: 50A×1D/A=50D

2. Additional Requirements from 100 B:

  • C: 100B×1C/B=100C

  • D: 100B×1D/B=100D

3. Total C and D After B Breakdown:

  • Total C: 150(from A)+100(from B)=250C

  • Total D: 50(from A)+100(from B)=150D

4. Additional Requirements from 250 C:

  • D: 250C×1D/C=250D

  • E: 250C×2E/C=500E

5. Total D After C Breakdown:

  • Total D: 150(previous)+250(from C)=400D

(Note: D requires F and G, but these do not affect E.)



Final Gross Requirement for E:

  • Only C directly demands E, and we’ve already calculated:

    250C×2E/C=500E


Summary:

  • E is only needed for C.

  • Total C required = 250 (from A and B).

  • Each C requires 2 E, so:

    250×2=500E

Answer: The gross requirement for component E is 500 units


answered by: anonymous